OpenStudy (anonymous):

derivative of (x^2+2)^2/3

OpenStudy (anonymous):

Hello, 123... You need to use the chain rule. If you let u=x^2+2, then\[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}=\frac{d}{du}(\frac{u^2}{3}).\frac{d}{dx}(x^2+2)\]\[=\frac{2u}{3}.2x\]\[=\frac{2}{3}(x^2+1)(2x)=\frac{4x(x^2+1)}{3}\]

OpenStudy (anonymous):

x^4+4 / 3